Please Do everything in redCARLA VISTA COMPANY Balance Sheets December 31 Assets 2019 2018 Cash $ 70,000 $ 67,000 Debt investments (short-term) 51,000 4(a1) Compute the liquidity and profitability ratios of Carla Vista Company for 2018 and 2019. (Round all answers to 2 decimal
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Carla Vista A B C=B-A D=C/A
Income Statement 2019 2018 Answer a 1 2018 2019 Change Change %
Sales 908,000.00      796,000.00 Current Ratio     2.0     1.7
Cost of goods sold 647,000.00      576,000.00 Current Assets/Current Liabilities
Gross Profit 261,000.00      220,000.00
Operating Expense 194,000.00      160,000.00 Earnings per share 2.86 3.19      0.33 11.67%
Net Income     67,000.00        60,000.00 Net Income/ Number of shares
Balance Sheet
Assets 2019 2018
Current Assets
Cash      70,000.00        67,000.00
Debt Investments (short term)      51,000.00        41,000.00
Accounts Receivable (net) 104,000.00        90,000.00
Merchandise Inventory 230,000.00      164,000.00
Prepaid Expenses      22,000.00        23,000.00
Total Current Assets 477,000.00      385,000.00
Land 133,000.00      133,000.00
Building and Equipment, net 262,000.00      185,000.00
Total Assets 872,000.00      703,000.00
Liabilities & Stockholders' Equity
Liabilities
Notes Payable 172,000.00      102,000.00
Accounts Payable      66,000.00        51,000.00
Accrued Liabilities      39,000.00        39,000.00
Current Liabilities 277,000.00      192,000.00
Bonds payable, due 2022 249,000.00      169,000.00
Total Liabilities 526,000.00      361,000.00
Common Stock $ 10 par 210,000.00      210,000.00
Retained Earnings 136,000.00      132,000.00
Total Stockholders' Equity 346,000.00      342,000.00
Total Liabilities & Stockholders' Equity 872,000.00      703,000.00
Common Stock $ 10 par 210,000.00      210,000.00
Number of shares     21,000.00        21,000.00
2019 2020 Note
Number of shares sold        19,000.00 D
Par value per share                10.00 E
Value of shares sold      190,000.00 F=D*E
Add: Net Income for 2020        52,000.00
Add: Stockholders' Equity for 2019      342,000.00
Stockholders' Equity 346,000.00      584,000.00 G
Net Income      67,000.00        52,000.00 H
Return on common stockholders equity 19.36% 8.90% I=H/G
Net Income/ Stockholders equity
Change % -10.46% This is the difference between 2020 and 2019 figure.
Total Liabilities in 2019      526,000.00
Less: Notes payable paid      172,000.00
Total Liabilities 526,000.00      354,000.00 J
Total Assets 872,000.00      864,000.00 K
Debt to Assets ratio 60.32% 40.97% L=J/K
Total Liabilities/ Total Assets
Change % -19.35% This is the difference between 2020 and 2019 figure.
Number of shares      21,000.00        40,000.00 For 2020, its 21,000 shares plus 19,000 further issued in 2020.
Net Income      67,000.00        52,000.00
Earnings per share                3.19                  1.30 M
Net Income/ Number of shares
Marker price per share                9.00                13.00 N
Price Earnings Ratio                  2.8                  10.0 O=N/M
Marker price per share/ Earnings per share
Change                  7.18 This is the difference between 2020 and 2019 figure.
Change % 254.50% This is the change/ 2019 figure.